OpenStudy (paradise):

The tour includes a trip to one museum, one amusement park, and one restaurant. There are four different museums, three amusement parks, and five restaurants to choose from. How many choices are there for your tour?

OpenStudy (austinl):

so you have 1/3 * 1/4 * 1/5 1/(3*4*5)

OpenStudy (austinl):

It shouldn't be divided. It is just 3*4*5. My apologies, I am still rusty with statistics.

OpenStudy (paradise):

\[\frac{ 1 }{ 3*4*5 }= \frac{ 1 }{ 60 }\]

OpenStudy (paradise):

so just 60?

OpenStudy (austinl):

It should be, yes, there are 60 different variations of it.

OpenStudy (paradise):

oh okay thank you(:

OpenStudy (austinl):

No problem!